# Fine-tuning with Very Large Dropout

### v2

@@ 4.2.1 Large learning rates for the last layer
[…]
4.2.2 Ensemble and weight averaging
Figure 5: Incremental benefits achieved by constructing ensembles or by averaging the weights of models fine-tuned with very large dropouts. The baseline and dropout results are the same as those reported in Figure 7. In contrast, the ensemble and weight averaging results are now obtained by averaging the output or the weights of models fine-tuned with large dropouts. Ensemble and weight averaging techniques provide a marginal o.o.d. performance improvement on vlcs or office home and a negligible o.o.d. performance improvement on pacs or terra incognita.
@@ 4.2.2 Ensemble and weight averaging
Figure 5 similarly explores the incremental benefits achieved by constructing ensembles or by averaging the weights of models fine-tuned with very large dropouts. The incremental improvements in o.o.d. performance achieved by these methods, if any, are much smaller than the improvement achieved by large dropout rates alone. Comparing Figures 7 and 5 also shows that in contrast, fine-tuning with large dropout rates before computing ensembles or averaging model weights brings large o.o.d. performance improvements over fine-tuning without dropout.
dataset ResNet #1 naive fine-tune ResNet #1 advance fine-tune ResNet #2 naive fine-tune ResNet #2 advance fine-tune
vlcs 76.7 79.5 78.3 81.0
office home 68.9 71.8 71.4 74.3
pacs 86.2 87.5 87.3 89.2
terra incognita 48.2 49.7 51.0 55.8
Average 69.0 72.1 72.0 75.1
@@
4.3 Richer pre-training beats sophisticated fine-tuning
We have demonstrated that the very-large dropout method delivers consistently better o.o.d. performance than computing ensembles or weight-averages of models fine-tuned without dropout. However we also have argued that fine-tuning does not create new representations but merely exploits the representations already present in the pre-trained model. Therefore the final o.o.d. performance of this fine-tuning process must strongly depend on the quality and the diversity of the features present in the pre-trained network (richer representation), even if these features are not exploited by the pre-trained network but buried in its hidden layers.
To validate this assertion, we compare the i.i.d. and o.o.d. performance obtained by various methods applied to ResNet50 networks pre-trained using the same ImageNet data but using different data augmentation schemes. As explained in the first paragraphs of section 4, the results reported so far use a network pre-trained using a broad array of data augmentation techniques, termed ResNet #2. We now compare its fine-tuning properties with network termed ResNet #1 pre-trained using the simpler protocol described in He et al. (2016).
Figure 3 compares the o.o.d. performances of both networks after regular fine-tuning and after fine-tuning with all the available tricks, that is, with dropout, with 10×10$\times$10 × larger last layer learning rate, and after averaging the weights of checkpoints collected along the fine-tuning trajectory. This comparison makes clear that the quality of the pre-trained representation matters more than the sophistication of the fine-tuning techniques. This is consistent with the idea that fine-tuning only leverages the existing features of the pre-trained network and does not create new ones.
@@ Appendix A Experiment details
A.1 Fine-tuning DomainNet in Figure 4
The DomainNet o.o.d. fine-tuning experiment in Figure 4 follows the same pipeline as other o.o.d. fine-tuning experiments on vlcs, pacs, Office Home, and Terra Incognita datasets. Due to the larger size of DomainNet dataset, we use larger learning rates $\{3.10^{-3},5.10^{-3}\}$ and a longer training iteration $30,000$ with a 10% learning rate decay at $15,000$ and $25,000$.
A.2 Training from scratch in Figure 4
@@ A.3 Compute Resources
Appendix B Additional results
B.1 Residual block dropout (stochastic depth) in o.o.d. fine-tuning
Huang et al. (2016) shows dropping a residual block at random with probability $p$ helps train very deep residual neural networks from scratch. Dropping a residual block, for example $f_{1}(x)$, in $\Phi(x)$ blinds the feature $\phi_{1}(x)$ but also changes the input of successive layers:
[…]
So an aggressive residual block dropout heavily disturbs the input of high residual blocks, hurts the fine-tuning process. Figure 6 showcase the effect of redisual block dropout in o.o.d. fine-tuning.
Figure 6: Residual block dropout (stochastic depth) and penultimate layer representation dropout comparison. Residual block dropout random drops a residual block with probability $\in\{0.1,0.2,0.3\}$. Mild residual block dropouts (0.1) provides some o.o.d. improvements (), but lags behind dropout () on vlcs, office home, and terra incognita. Aggressive residual block dropouts hurt fine-tuning (e.g. $\geq 0.3$ on office home and terra incognita), because dropping a residual block change the input of successive layers. Likewise the comparison of ensemble, weight averaging, and large last-layer learning rate in section 4.2, a proper residual block dropout helps fine-tuning in general, but is still a secondary factor compared with dropout.
B.2 o.o.d. performance improvement is not a secondary effect of some i.i.d. performance improvement
Figure 7 provides both i.i.d. and o.o.d. performance of very large dropout, ensembles and weight averaging methods. The very large dropout method outperforms other methods on o.o.d. performance without any advantage on i.i.d. performance. It reveals that the o.o.d. performance gaps of different methods in Figure 7 and Figure 1 do not come from the i.i.d. overfitting/underfitting.
[…]

### v3

@@ 4.2.1 Very-large dropout
4.2.2 Very-large dropout
$+$ ensemble or weight averaging
@@ 4.2.2 Very-large dropout
Table 3 similarly explores the incremental benefits achieved by constructing ensembles or by averaging the weights of models fine-tuned with very large dropouts. The results show that very-large dropout approach is compatible with ensembles and weight averaging apporach to gain a non-negative incremental imporvements in o.o.d. performance. On the other hand, comparing Table 1 and 3 shows that fine-tuning with large dropout rates before computing ensembles or averaging model weights brings large o.o.d. performance improvements over fine-tuning without dropout.
In short, the very-large dropout approach is compatible with other fine-tuning techniques but acts as the leading factor in terms of o.o.d. performance.
4.3 Robustness to hyperparameter selection
Out-of-distribution finetuning performance is known to be sensitive to hyperparameter selection (Ahuja et al., 2020; Wortsman et al., 2022a). To reduce the uncertain of hyperparameter selection, Figure 1 presents the box plot of different hyperparameter combinations (where each choice of training sub-datasets searches 6 hyperparameter combinations).
On all four datasets, the bottom of very-large dropout box (25% quartile) outperforms the top of other baseline boxes (75% quartile). On office home and pacs datasets, there is even a large gap between the worst dropout results and the best baseline results.
Figure 1: o.o.d. performance comparison between very large dropout, ensembles, and weight averaging methods on four DomainBed tasks. erm results were obtained using plain fine-tuning with different hyperparameters. Weight averaging results either average the model weights collected every 300 iterations along each fine-tuning trajectory or the final model weights of all fine-tuning trajectories as in (Ramé et al., 2022b). Ensemble results average instead the model outputs. Finally, large dropout results were obtained like the erm results but using a 90% dropout rate on the penultimate layer. Each box summarizes the results obtained with different hyper-parameters combinations.
4.4 Robustness of dropout rate selection
To the best of our knowledge, such large dropout rates (90% and above) are considered unsuitable for training a network from scratch and have not been previously used for fine-tuning either. This section study the relationship between dropout rates and o.o.d. performance. A smooth relationship indicates the robustness of dropout rate selection, while a curly relationship reflects the sensitivity.
Table 4 compares various dropout rates on the four tasks. A $90\%$ dropout rate reliably produces good o.o.d. performance on all four tasks. The optimal dropout rate for o.o.d. performance ranges from 90% to 95% for vlcs and pacs task (with 10k examples). And becomes slightly smaller, about 90%, for the slighlty larger datasets office home and terra incognita (with 15k to 25k examples).
Furthermore, the relationship between dropout rate and o.o.d. performance are smooth on all four datasets, which makes it easy to select the right dropout rate.
Table 4: Effect of diverse dropout rates during fine-tuning. The best o.o.d. performances are attained using rates around or above 90%. A large dropout rate (e.g. 90%) reliably produces good o.o.d. performance on all four tasks.
dropout rate 0% 50% 90% 95%
vlcs 78.3 79.7 80.1 80.4
office home 71.4 73.1 73.6 73.0
pacs 87.3 88.0 88.5 88.4
terra incognita 51.0 52.4 53.9 52.3
4.5 When should one apply very-large dropout?
We have demonstrated that the very-large dropout method delivers consistently better o.o.d. performance than computing ensembles or weight-averages of models fine-tuned without dropout. However we also have argued that fine-tuning does not create new representations but merely exploits the representations already present in the pre-trained model. Therefore the final o.o.d. performance of this fine-tuning process must strongly depend on the quality and the diversity of the features present in the pre-trained network (richer representation), even if these features are not exploited by the pre-trained network but buried in its hidden layers. i.e. the scope of applying very-large dropout method lies in situations where a rich representation has already been established.
Of course, modern foundational models, where many features are learned from a large and carefully constructed dataset, make this condition relatively easy to achieve. Thus provide a large space to apply this very-large dropout approach.
In this section, we study this condition precisely. We first study the performance of very-large dropout approach on the scratch-training scenario, where the representation is random. Then we progressively enrich the representation by pretraining and pretraining with enormous augmentations.
Random initialization and representation.
Figure 2 shows the effect of various dropout rates when one trains a network on the vlcs task from scratch, that is starting from a randomly initialized network without pretraining (i.e. random initialization and random representation). The optimal dropout rate falls to about zero. Dropout rates higher than 50% have a negative impact on both the i.i.d. and the o.o.d. performance of the network. This suggests that high dropout rates make it difficult to create new features (a nonlinear operation), but does not prevent leveraging existing features that were possibly buried in the network inner layers (a linear operation). This is the idea of richer representation we discussed in section 1.
Richer and richer representation.
@@ 4.5 When should one apply very-large dropout?
dataset ResNet #1 erm ResNet #1 very-large dropout ResNet #2 erm ResNet #2 very-large dropout
vlcs 76.7 78.1 78.3 80.1
office home 68.9 69.1 71.4 73.6
pacs 86.2 86.5 87.3 88.5
terra incognita 48.2 48.8 51.0 53.9
Average 70.0 70.6 72.0 74.0
To study the impact of rich representation, we compare the o.o.d. performance obtained by various methods applied to ResNet50 networks pre-trained using the same ImageNet data but using different data augmentation schemes. As explained in the first paragraphs of section 4, the results reported so far use a network pre-trained using a broad array of data augmentation techniques, termed ResNet #2. We now compare its fine-tuning properties with network termed ResNet #1 pre-trained using the simpler protocol described in He et al. (2016).
Table 5 compares the o.o.d. performances of both networks after regular fine-tuning and after fine-tuning with very-large dropout. Note that ResNet #2 contains richer representations than ResNet #1 due to the vast data augmentations. On ResNet #1, where the representation is richer than random representation, a very-large dropout rate (0.9) starts to help o.o.d. performance (0.6%). On ResNet #2, where the representation is richer than ResNet #1, the same very-large dropout approach vastly boosts o.o.d. performance (2%).
The results in this section showcase an increasing o.o.d. benefits of the very-large dropout approach as the representation getting richer. Starting from the scale of ResNet50 and ImageNet, the o.o.d. benefits of a very large dropout becomes significant.
In the context of large foundational models, both model size and dataset size are far larger than ResNet50 neural network and ImageNet dataset. Thus the space to apply this very-large dropout approach is large.
@@ 5 Discussion
Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Experiment details
A.1 Training from scratch in Figure 2